OpenStudy (anonymous):

what is the slope using the pairs (-3,4) and (2,-3)

OpenStudy (anonymous):

\[-\frac{7}{5}\]

OpenStudy (anonymous):

5/-7= -5/7 i think been a while

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

7/-5

OpenStudy (anonymous):

the y's go on top

OpenStudy (anonymous):

Rise/Run=(y2-y1)/(x2-x1)=-3-4/2-(-3)=-7/5